OpenStudy (anonymous):

Describe the distributive property? How is it different from other properties? Can you give an example?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

x(2x + 3) Using the distributive property, you can simplify: x(2x + 3) = (x)(2x) + (x)(3) = 2x^2 + 3x

OpenStudy (anonymous):

@StudyGurl14 can you describe what distributive property is first please? :)

OpenStudy (studygurl14):

I don't really know how to explain it, sorry. It's just one of those things that are better explained with an example